OpenStudy (anonymous):

You have 750,000 IRA at time you retire. You have the option of investing this money in two funds. Funds A pays 1.3% annually and Fund B pays 6.3% annually. How should you divide your money between A and B to produce an annual interest income of $37,750

OpenStudy (stamp):

Sounds like a system of equations, that is my opinion at a glance, I have not delved any further.

OpenStudy (anonymous):

thats what i thought but cant find the steps to complete it in my textbook

OpenStudy (anonymous):

Is this simple or compound interest?

OpenStudy (anonymous):

to be honest im not even sure i got for fund A- 560000 and fund B- 190000

OpenStudy (anonymous):

Then let's assume simple. I haven't solved for anything on my own yet, but here's the setup so far. Let's say you take x amount of dollars from the total $750,000 and put it into Fund A. So, you're left with (750,000-x) to put into Fund B. The balances of Funds A and B will be given by A_1 and A_2, respectively. \[A_1=P_1(1+r_1t)\\A_2=P_2(1+r_2t),\] where A is the balance after time t, P = initial balance/deposit, r = (annual simple) interest rate, and t = time (which we'll leave as 1 year). So, you have \[A_1=x(1+0.013)\\ A_2=(750,000-x)(1+0.063)\] Does the setup make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, after one year, the balance of the accounts is \[\begin{align*}A_1+A_2&=1.013x+1.063(750,000-x)\\ &=797,250-0.05x\end{align*}\]

OpenStudy (anonymous):

Since you're working with simple interest, you'll be using the formula \[I=Prt\] In this case, P = A_1 + A_2 and t = 1. One minor change, though. A_1 and A_2 have different rates associated with each other, so you'll have to write the interest equation as \[I=(A_1r_1+A_2r_2)\] Do you follow?